OpenStudy (anonymous):

Which of the following is a factor of 27x3 – 64y3?

OpenStudy (tkhunny):

3x - 4y

OpenStudy (anonymous):

You have the difference of 2 perfect cubes and that can be factored as follows: a^3 - b^3 = (a - b)(a^2 + ab + b^2) where a = 3x and b = 4y

OpenStudy (anonymous):

I'm pretty sure this is the same question on an assignment I'm doing. I think one of the factors is 9x^2+12xy+16y^2 Hope this helps!